A computer is infected with the Sasser virus. Assume that it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another five minutes, etc. How long until 100 million computers are infected?
Answer by scott8148(6628)

let x="time in minutes" __ 20^(x/5)=100000000

(x/5)log(20)=log(100000000) __ x/5=8/log(20) __ x=40/log(20)